OpenStudy (anonymous):

Which method of integration do you need to use to integrate: 9ln(x)/x^2 ?

OpenStudy (loser66):

substitute

OpenStudy (anonymous):

sometimes it's just staring you right in the eye ^.^

OpenStudy (anonymous):

subsitute what? if we let u = ln (x), du = 1/x

OpenStudy (anonymous):

yeah? And what's wrong with that? :D

OpenStudy (anonymous):

your integral will be in terms of u and x not u only. You can't integrate it.

OpenStudy (anonymous):

oh jeez :D If we let u = ln(x) \[\Large \int \frac{9\ln(x)}{x^2}dx \rightarrow \int \frac{9u}{x}du\rightarrow \int \frac{9u}{e^u}du\] Happy? :D

OpenStudy (anonymous):

\[\LARGE \rightarrow \int 9ue^{-u}du\] Let's do this by parts... when we feel like it ^.^ ta

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

Hold on A SEC

OpenStudy (anonymous):

\[9\int\limits_{}^{}lnx*\frac{ 1 }{ x^2 }=-9\int\limits_{}^{}lnx*(\frac{ 1 }{ x })'=-9[lnx*\frac{ 1 }{ x}-\int\limits_{}^{}\frac{ 1 }{ x^2 }]\] \[=-9[\frac{ lnx+1 }{ x}]+\zeta\]

OpenStudy (anonymous):

I've used the integration by parts